OpenStudy (anonymous):

Simplify by removing factors of 1. q^2-49/q^2-14q+49

jhonyy9 (jhonyy9):

(q-7)(q+7) ---------- = (q+7)/(q-7) (q-7)(q-7)

OpenStudy (aroub):

= (q-7) (q+7) ----------- (q-7) (q-7) = q+7 ------ q-7 \[q \neq 7 \]